# Rheo-SINDy: Finding a constitutive model from rheological data for complex fluids using sparse identification for nonlinear dynamics

### v2

@@ V Concluding Remarks
Rheo-SINDy for two phenomenological constitutive equations (i.e., the upper convected Maxwell model and Giesekus model) revealed the following two things. First, compared to constant shear tests, oscillatory shear tests are appropriate for generating training data. Second, the sequentially thresholded Ridge regression (STRidge) and adaptive Lasso (a-Lasso) are effective in finding appropriate constitutive equations. We then examined the commonly used mesoscopic model, namely the dumbbell model with three different representations of spring strength: the Hookean, FENE-P, and FENE springs. Although the Hookean and FENE-P dumbbell models have analytical constitutive equations, for the FENE dumbbell model, there is no analytical expression of the constitutive equation. We confirmed through the Hookean dumbbell model that even in the presence of noise, the a-Lasso provides the correct solution over a wide range of the hyperparameter $\alpha$. Rheo-SINDy was also effective in discovering the complex constitutive equations for the FENE-P dumbbell model. This case study revealed that the identification of complex equations requires the preparation of an appropriate custom library based on prior physical knowledge. Using physical insights obtained from the Hookean and FENE-P dumbbell models, we attempted to find approximate constitutive equations for the FENE dumbbell model. We found that the a-Lasso can successfully give the approximate constitutive equations, which can be used in predictions beyond the range of the training data.
@@ Acknowledgements.
We thank Prof. Yoshinobu Kawahara for his valuable comments on data-driven methods. We also thank Dr. John J. Molina for carefully reading the manuscript and providing insightful comments. TS was partially supported by JST PRESTO Grant Number JPMJPR22O3. SM was financially supported by the Kyoto University Science and Technology Innovation Creation Fellowship (FSMAT), Grant Number JPMJFS2123.

### v3

@@ V Concluding Remarks
Rheo-SINDy was applied for two phenomenological constitutive equations (i.e., the upper convected Maxwell model and Giesekus model), which revealed the following two things. First, compared to constant shear tests, oscillatory shear tests are appropriate for generating training data. Second, the sequentially thresholded Ridge regression (STRidge) and adaptive Lasso (a-Lasso) are effective in finding appropriate constitutive equations. We then examined the commonly used mesoscopic model, namely the dumbbell model with three different representations of spring strength: the Hookean, FENE-P, and FENE springs. Although the Hookean and FENE-P dumbbell models have analytical constitutive equations, for the FENE dumbbell model, there is no analytical expression of the constitutive equation. We confirmed through the Hookean dumbbell model that even in the presence of noise, the a-Lasso provides the correct solution over a wide range of the hyperparameter $\alpha$. Rheo-SINDy was also effective in discovering the complex constitutive equations for the FENE-P dumbbell model. This case study revealed that the identification of complex equations requires the preparation of an appropriate custom library based on prior physical knowledge. Using physical insights obtained from the Hookean and FENE-P dumbbell models, we attempted to find approximate constitutive equations for the FENE dumbbell model. We found that the a-Lasso can successfully give the approximate constitutive equations as combinations of known elemental terms given in a function library, which can be used in predictions beyond the range of the training data. Finally, we present a regression method that yields a constitutive model satisfying rotational indifference. These results indicate that Rheo-SINDy is ready to be applied to data obtained from more complex models or experimental data.
@@ Acknowledgements.
We thank Prof. Yoshinobu Kawahara for his valuable comments on data-driven methods. We also thank Dr. John J. Molina for carefully reading the manuscript and providing insightful comments. We are grateful to Prof. Takashi Taniguchi for encouraging us to start this research. TS was partially supported by JST PRESTO Grant Number JPMJPR22O3. SM was financially supported by the Kyoto University Science and Technology Innovation Creation Fellowship (FSMAT), Grant Number JPMJFS2123.